How many double bonds are present in C10H18? This compound consumes 2 mol of H2 on catalytic hydrogenation.
Answer #1

C10H18 molecule has double bond equivalent of

DBE = C+1 - (H/2)

DBE = 10 + 1 - (18/2)

= 11- 9 =2

Thus the degree of unsaturation is 2

As it is already given that 2 mole of H2 is required for the catalytic hydrogenation of the compound.

Since 1 mole of H2 is required to reduce a single doule bond. Similarly 2 moles are required for 2 double bonds.

ANSWER:

number of double bonds =2